OpenStudy (anonymous):

(Multiple Choice Worth 1 points) open parentheses 5 plus square root of 3 close parentheses open parentheses 5 minus square root of 3 close parentheses 22 28 22 minus10 square root of 3 25 minus 12 square root of 3

OpenStudy (anonymous):

\[(5+\sqrt{3})(5-\sqrt{3})=22\]

OpenStudy (anonymous):

it says there was a math processing error?

OpenStudy (anonymous):

(5 + sqrt 3)(5 - sqrt 3) = 5^2 - sqrt 3^2 = 25 - 3 = 22 (a + b)(a - b) = a^2 - b^2

OpenStudy (anonymous):

What kind of program are you using? If it is multiple choice, and you post the right question, then 22 is the right answer